# USTC-ICTS/PCFT-25-37

### v1

@@ I Introduction
Loop-level Feynman diagrams play a central role in perturbative quantum field theories (QFTs). In the era of precision physics, The current era of precision physics presents an urgent demand for the calculation of multi-loop amplitudes, which are crucial for collider phenomenology, gravitational wave predictions from black hole scattering, as well as formal aspects of QFT. In multi-loop computations, a key step is to reduce all Feynman integrals to a finite, linearly independent set consisting of master integrals. This step, known as Feynman integral reduction, is frequently the computationally intensive bottleneck of perturbative QFT calculation. Conventionally, integral reduction is carried out using the integration-by-parts (IBP) identities [1, 2] via the linear reduction algorithm named after Laporta [3]. However, for most cutting-edge multi-loop computations, the IBP method inevitably involves a vast linear system, which is prohibitively difficult to solve via linear reduction.
@@
The main bottleneck of integration-by-parts (IBP) reduction arises in multi-loop integrals with high-degree numerators or high-power propagators, where the number of IBP relations grows exponentially. This issue becomes particularly severe in recent cutting-edge multi-loop computations, such as the $5$ pm $1$ SF black-hole scattering [33] and the three-loop five-point Feynman integrals [34]. Tackling reduction problems at the three- and four-loop orders, or for two-loop multi-leg processes, necessitates new conceptual breakthroughs.
Recurrence rules [35, 9] provide an efficient alternative to IBP reductions, particularly for the challenging task of reducing multi-loop integrals with high-rank numerators or elevated propagator powers. The implements of recurrence rules include methods based on gröbner basis [36, 37, 35, 38, 39] and heuristic algorithms [9, 10]. However, a significant drawback was the computational cost associated with the recurrence rule derivation, which has limited their widespread adoption.
In this work, we present a novel and highly efficient method, based on generating functions [40, 41], to systematically derive complete symbolic recurrence relations for the reduction of Feynman integrals. By expressing Feynman integrals in terms of generating functions, integration-by-parts (IBP) relations are translated into differential equations for these generating functions. The reduction of Feynman integrals is then transformed into the task of reducing a differential operator with respect to these differential equations. We introduce a new approach for differential operator reduction, based on operator grading and integrability conditions, which allows us to efficiently tobtain symbolic reduction rules for arbitrary Feynman integrals.
This paper is organized as follows. We first review the definition of generating functions for Feynman integral reduction. We then focus on our new systematic method for obtaining symbolic reduction rules. As illustrations, we apply this method to three multi-loop reduction examples: the sunset diagram, and planar and non-planar double-box diagrams.
@@ II Generating Functions for Feynman integrals
For $L$-loop integrals with $E$ independent external momenta, a family of scalar FIs can be represented as
@@
Since the generating function encodes all Feynman integrals in the family, it is sufficient to construct IBP relations for the generating function alone, which we call as the fundamental IBP relations. These relations yield DEs of the form:
@@
where $B$ is the function of generating functions of subsectors and $c_{t}$ are coefficients. We define the degree of an operator as $\sum_{i=1}^{N}(\vec{o})_{i}$, where the index vector is $\vec{o}\equiv\vec{b}-\vec{a}$. By substituting the expansion (3) to equation (5) and extracting the coefficient of $\vec{\eta}^{\vec{n}}$, we obtain
@@
where for simplicity, we have made the convention that $F_{\vec{n}}=0$ if any component $(\vec{n})_{i}<0$. We call equation (6) as the recurrence relation derived from the DE (5).
Within this framework, the goal of symbolic reduction is transformed into an algebraic problem: expressing a higher-degree operator as a linear combination of lower-degree operators, using the DEs derived from the fundamental IBP relations. The primary objective of this work is to develop a systematic algorithm to achieve this task.
@@ III Symbolic Reduction Rules
The DEs derived from fundamental IBP identities can be classified based on the highest-degree operators. This classification directly informs how symbolic reduction rules are extracted:
  • I:
    The equation contains a single, unique highest degree operator, which can be directly solved to yield an immediate reduction rule.
  • II:
    The equation contains multiple highest degree operators that share the same operator index. When translated into the recurrence form (6), they effectively leave a single highest-degree term, yielding a reduction rule.
  • III:
    The equation contains multiple highest degree operators with different operator indices. In this case, one operator is chosen as the target to solve, according to the aim of reduction.
Given a reduction rule, i.e, expressing a higher operator index $\vec{o}$ by others, we can determine which points can be reduced or can not. They are given by two sets, the reduced set
${\cal S}_{\widehat{O}}=\{\vec{n}\in Z^{N}|\vec{n}\geq 0\penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ \vec{n}-\vec{o}\geq 0\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(7)
and unreduced set
${\cal U}_{\widehat{O}}=\{\vec{n}\in Z^{N}|\vec{n}\geq 0\penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ \exists i,(\vec{n}-\vec{o})_{i}<0\}\penalty 10000\ .\penalty 10000\ \penalty 10000\ \penalty 10000\$(8)
A crucial component of our methodology is the generation of "descendant" rules by acting on an existing DE with differential operators. The complete algorithm is structured as an iterative loop that expands the set of known reduction rules until a complete basis is found.
We now outline the core algorithm for deriving a complete set of symbolic reduction rules:
  1. (A)
    Initialization: Generate an initial set of DEs from the fundamental IBP relations. Extract immediate reduction rules and define the initial reduced set $\mathcal{S}$ and unreduced set $\mathcal{U}$. Initialize the maximum degree of differential operators to $M=2$.
  2. (B)
    Generation of New Equations: Create a new set of DEs by applying differential operators $\prod_{i}(\partial/\partial\eta_{i})^{k_{i}}$, up to the current maximum degree $M$, to the existing DEs. Apply all known reduction rules (as well as their descendant reduction rules) to new DEs to eliminate as many as possible highest-degree operators in favor of lower-degree ones, which typically refers to integrability conditions of differential operators. In this way, we get the DE for new unsolved highest operators.
  3. (C)
    Finding new reduction rules: Using the obtained DE to solve new highest operators based on the current unreduced set $\mathcal{U}$. Simplifying new reduction rules by setting terms in DE’s to zero when applying these reduction rules to the set $\mathcal{U}$.
  4. (D)
    Update and Completeness Check: Based on new found reduction rules, update the reduced set $\mathcal{S}$ and unreduced set $\mathcal{U}$. Check if the size of the updated $\mathcal{U}$ matches the number of master integrals. If so, the reduction is complete and the algorithm terminates successfully. Else, return to step (B).
  5. (E)
    Expand the maximum degree of differential operators: If steps above failed to produce any new reduction rules, the system has stagnated at the current order $M$. Increment the maximum degree $M\to M+1$, and return to step (B) to generate more complex new equations and find new rules.
@@
Our framework offers distinct advantages over existing methods. Compared to methods based on gröbner basis [36, 37, 35, 38, 39], our approach circumvents complex non-commutative algebra. Instead, it systematically reduces the derivation of reduction rules to a more tractable linear algebra problem. Compared to heuristic algorithms [9, 10], our approach provides a general and reusable solution with a clear termination condition. This contrasts sharply with the ad-hoc nature of heuristic strategies, which lack guarantees of completeness and often fail when faced with exceptionally complex topologies.
@@ IV Example 1: the top sector of sunset diagrams
There are six fundamental IBP relations for top-sector, producing six type-I equations for degree-2 operators ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{j}}$, $i=4,5;j=1,2,3$ (see (22) to (27) for their explicit expressions). For example, the IBP relation $0=\int\prod_{i=1}^{L}\frac{\mathrm{d}^{D}\ell_{i}}{\mathrm{i}\pi^{D/2}}{\partial\over\partial\ell_{1}}\cdot\left\{\ell_{1}e^{\sum_{j=4,5}\eta_{j}s_{0}^{-1}\mathcal{D}_{j}}\prod_{i=1}^{3}{1\over\mathcal{D}_{i}-\eta_{i}s_{0}}\right\}$ produces
$0$ $=$ $\left\{2{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{3}}G_{111}\right\}-\left\{{2m_{1}^{2}\over s_{0}}{\partial\over\partial\eta_{1}}G_{111}+{f_{+-+}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}\right\}$ (11)

### v2

@@ I Introduction
Loop-level Feynman diagrams play a central role in perturbative quantum field theories (QFTs). The current era of precision physics presents an urgent demand for the calculation of multi-loop amplitudes, which are crucial for collider phenomenology, gravitational wave predictions from black hole scattering, as well as formal aspects of QFT. In multi-loop computations, a key step is to reduce all Feynman integrals to a finite, linearly independent set consisting of master integrals. This step, known as Feynman integral reduction, is frequently the computationally intensive bottleneck of perturbative QFT calculation. Conventionally, integral reduction is carried out using the integration-by-parts (IBP) identities [1, 2] via the linear reduction algorithm named after Laporta [3]. However, for most cutting-edge multi-loop computations, the IBP method inevitably involves a vast linear system, which is prohibitively difficult to solve via linear reduction.
@@
The main bottleneck of integration-by-parts (IBP) reduction arises in multi-loop integrals with high-degree numerators or high-power propagators, where the number of IBP relations grows exponentially. This issue becomes particularly severe in recent cutting-edge multi-loop computations, such as the conservative $5$ pm $2$ SF black-hole scattering [33, 34] which required $\sim 3\times 10^{6}$ core CPU hours with Kira3 [18], and the three-loop five-point Feynman integrals [35, 36] where the most complicated family’s IBP reduction used hundreds of unitarity cuts and tens of thousands IBP relations for each cut. Tackling reduction problems at the three- and four-loop orders, or for two-loop multi-leg processes, necessitates new conceptual breakthroughs.
Recurrence rules [37, 9] provide an efficient alternative to IBP reductions, particularly for the challenging task of reducing multi-loop integrals with high-rank numerators or elevated propagator powers. The implementations of recurrence rules include methods based on Gröbner bases [38, 39, 37, 40, 41] and heuristic algorithms [9, 10]. However, a significant drawback was the computational cost associated with the recurrence rule derivation, which has limited their widespread adoption.
In this work, we present a novel and highly efficient method, based on generating functions [42, 43], to systematically derive complete symbolic recurrence relations for the reduction of Feynman integrals. By expressing Feynman integrals in terms of generating functions, integration-by-parts (IBP) relations are translated into differential equations for these generating functions. Using IBP relations and other conditions (such as integrability conditions), the reduction of Feynman integrals is then reformulated as reducing differential operators based on operator grading. This method will enable us to efficiently obtain complete symbolic reduction rules for arbitrary Feynman integrals with arbitrary indices.
This paper is organized as follows. We first review the definition of generating functions for Feynman integral reduction. We then present our new systematic algorithm for obtaining symbolic reduction rules. As illustrations, we apply this method to three multi-loop reduction examples: the sunset diagram, the planar double-box diagram, and the non-planar double-box diagram.
@@ II Generating Functions for Feynman integrals
A Feynman integral family with $L$ loops and $E$ independent legs can be represented as
@@
The differential equations for generating function have the form
@@
where $B$ is the function of generating functions of subsectors, $\hat{O}_{t}$’s are the operator writing into the standard form and $c_{t}$ are coefficients. In the standard form, the operator is uniquely fixed by a pair of vectors $(\vec{a},\vec{b})$, which is called the finer index of operator. By extracting the coefficient of $\vec{\eta}^{\vec{n}}$, we obtain
@@
[…]
From (6), one can see the role of operators in the recurrence relation is largely fixed by the combination $\vec{o}\equiv\vec{b}-\vec{a}$, which is called the index of the operator. Similarly, we define the degree of an operator as $|\vec{o}|\equiv\sum_{i=1}^{N}(\vec{o})_{i}$, which relates to the degree of lattice points via (6). Furthermore, we call an equation of degree $M$, if it contains operators with the highest degree $M$. For an operator with more than one terms of the highest degree, the unique leading term would be defined in the next section for the Module II.
The mapping between (5) and (6) enables the reduction of Feynman integrals, i.e., obtaining coefficients of higher-degree lattice points in (3), becomes a problem to reduce a higher-degree operator as a linear combination of lower-degree operators using the DEs. The primary objective of this work is to develop a systematic algorithm to solve this task.
In the classic work [37], IBP reduction was reformulated as an operator reduction problem and solved via Gröbner basis. In this work, we emphasize that Gröbner basis is not used for reducing operators.
We comment on the operators. First, there are special operators with finer index $(\vec{a},\vec{a})$. which are called zero-index operators since $\vec{o}=\vec{0}$. [^2] Later, a zero-index operator could also mean a linear combination of zero-index operators. An important fact is that any two zero-index operators commute. Second, there are many operators with the same index. All of them can be written into the form $\widehat{O}_{0}\widehat{O}^{F}_{\vec{o}}$, where $\widehat{O}_{0}$ is a zero-index operator and $\widehat{O}^{F}_{\vec{o}}\equiv\vec{\eta}^{\vec{o}_{-}}{\partial^{\vec{o}_{+}}\over\partial\vec{\eta}^{\vec{o}_{+}}}$. The $\vec{o}_{\pm}$ are the positive and negative components of $\vec{o}$ such that $\vec{o}=\vec{o}_{+}-\vec{o}_{-}$. Thus, we introduce a useful concept: an operator $\widehat{Q}_{b}$ is called the descendant of an operator $\widehat{Q}_{a}$ if there is another operator $\widehat{O}_{1}$, such that
$\widehat{O}_{0}\widehat{Q}_{b}=\widehat{O}_{1}\widehat{Q}_{a},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(7)
where $\widehat{O}_{0}$ is a zero-index operator.
@@ III Symbolic Reduction Rules
The DEs can be categorized as the so-called T1-type or T2-type. An equation is the T1-type if all highest degree operators therein have the same operator index. Otherwise it is a T2-type equation. We would reduce the highest operators having a given index $\vec{o}$.
For a relation to be applicable for all choices $\vec{n}\geq 0$ in (6), it is required that $\vec{o}\geq 0$. Depending whether there is one or multiple operators for the index $\vec{o}$, the reduction can be divided into the A-type or B-Type. More explicitly,
$\text{T1A-type}:$ ${\partial^{\vec{b}}\over\partial\eta^{\vec{b}}}G_{\vec{\mu}}(\eta)=R_{\vec{b}}+B\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(8)
$\text{T1B-type}:$ $\widehat{O}_{0}{\partial^{\vec{b}}\over\partial\eta^{\vec{b}}}G_{\vec{\mu}}(\eta)=R_{\vec{b}}+B\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(9)
$\text{T2A-type}:$ ${\partial^{\vec{b}}\over\partial\eta^{\vec{b}}}G_{\vec{\mu}}(\eta)=Q_{\vec{b}}+R_{\vec{b}}+B\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(10)
$\text{T2B-type}:$ $\widehat{O}_{0}{\partial^{\vec{b}}\over\partial\eta^{\vec{b}}}=Q_{\vec{b}}+R_{\vec{b}}+B\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(11)
where $\widehat{O}_{0}=\left[c_{0}+\sum_{i}c_{i}\eta^{\vec{a}_{i}}{\partial^{\vec{a}_{i}}\over\partial\eta^{\vec{a}_{i}}}\right]$ is a zero-index operator. The $Q$ part contains operators with the same degree as the left hand side, while the $R$ part contains only operators with lower degree. Finally $B$ contains contributions from sub-sectors. Importantly, we further require $\widehat{O}_{0}$ containing identity operator for (9) and (11) (i.e., $c_{0}\neq 0$).
Our novel algorithm can be summarized by the diagram with three modules, or a golden triangle,
I: Generating EquationsII: Solving Equations to Get Reduction RulesIII: Reduction rules Complete?fundamental IBPsFinishYesNo
We present a brief explanation for each module:
  • Module I: Input DEs come from two different origins: (a) the initial seeds of DEs from fundamental IBP relations; (b) all others from the action of $\partial\over\partial\eta_{i}$} on reduction rules.
    After the action of $\partial\over\partial\eta_{i}$}, a crucial step is to use all known reduction rules to simplify it. For the highest degree operator $\widehat{Q}_{b}$, if it is the descendant of the reduction rule $\widehat{Q}_{a}$ then in the following combination according to (7),
    $\widehat{Q}_{0}{\rm Eq}-\widehat{Q}_{1}{\rm Rule}_{\widehat{Q}_{a}}$ (12)
    eliminate the highest degree part.
    Repeat this procedure until there is no highest degree operators to be simplified. If all highest operators have been reduced, we move to the next highest degree operators and repeat.
  • Module II: Classify equations from the module I as different subsets according to degrees. Then we start solving the set with highest degree and then deal with lower-degree sets, until we reach the degree zero. For each subset, the highest degree operators can be reduced by linear algebra methods, say, Gaussian elimination.
    When solving the T2-type equations, which have highest degree operators with different operator indices, we must carefully define a global priority ordering. The core policy is to prevent the formation of loops in the reduction rules, which is achieved by ensuring a net decrease in at least one index throughout the recursive process, which guarantees the eventual termination of the reduction chain.
    After solving the subset, new reduction rules are implemented to simplify previous reduction rules and remaining unsolved equations. This step is called update.
  • Module III: With a reduction rule for the operator with index $\vec{o}$, the lattice points that can be reduced by this rule are in the set,
    ${\cal S}_{\vec{o}}$ $=$ $\{\vec{m}\in\mathbb{Z}^{N}|\vec{m}\geq 0\penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ \vec{m}-\vec{o}\geq 0\}\,$$penalty 10000\ \penalty 10000\ \penalty 10000\$(13)
    and the irreducible lattice points are in the set,
    ${\cal U}_{\vec{o}}=\{\vec{m}\in\mathbb{Z}^{N}|\vec{m}\geq 0\penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ \vec{m}-\vec{o}\not\geq 0\}\,$$penalty 10000\ \penalty 10000\ \penalty 10000\$(14)
    Thus, with all reduction rules, the set of irreducible lattice points at this point is ${\cal U}_{total}\equiv\cap_{i=1}^{P}{\cal U}_{\vec{o}_{i}}$. Then count the number of lattice points in the set. if the number equals the number of master integrals in this sector, reduction rules are complete. Otherwise, start a new round of computations with three modules.
@@
Our framework offers distinct advantages over existing methods. Compared to methods based on Gröbner bases [38, 39, 37, 40, 41], our approach circumvents complex non-commutative algebra. Instead, it systematically reduces the derivation of reduction rules to a more tractable linear algebra problem. Compared to heuristic algorithms [9, 10], our approach provides a general and reusable solution with a clear termination condition. This contrasts sharply with the ad-hoc nature of heuristic strategies, which lack guarantees of completeness and often fail when faced with exceptionally complex topologies.
@@ IV Example 1: the top sector of sunset diagrams
This is a simplest example to clearly show our algorithm. The sunset diagram has the propagators and ISPs,
$\mathcal{D}_{3}$ $=$ $((\ell_{1}+\ell_{2}-K)^{2}-m_{3}^{2}),$
The details of computations can be found in the Supplemental Material, here we will just present the outline of the whole procedure.
The first round of reduction: The initial seed for top-sector $G_{11100}=G_{111}$ is six degree two DE’s from fundamental IBP relations. Using them, we find $6$ T1A-type reduction rules for degree-2 operators. ${\partial\over\partial\eta_{a}}{\partial\over\partial\eta_{i}}$, $a=4,5;i=1,2,3$. Using them the irreducible lattice points are $(0,0,0,n_{4},n_{5})$ and $(n_{1},n_{2},n_{3},0,0)$.
The second round of reduction: Now generate new equations by applying $\partial\over\partial\eta_{i}$} over the $6$ reduction rules above. Then we get $9$ nontrivial DE’s, out of which only $8$ are independent. The degree of the DE’s depends on the mass configuration.
For the massless case, there are $3$ degree-two DE’s and $5$ degree-one DE’s. Using them, we can find T1B-type reduction rules and T2B-type reduction rules for Thus the irreducible lattice points are just $(0,0,n_{3},0,0)$.
For the massive case, there are $5$ degree-two DE’s and $3$ degree-one DE’s. Using them, we find T2B-type reduction rules for and $5$ T2A-type reduction rules. The irreducible points consist of the set $(0,0,n_{3},0,0)$ and points $(1,0,0,0,0)$, $(0,1,0,0,0)$.
The third round of computation: Now we act $\partial\over\partial\eta_{3}$} on the reduction rule of, for example, $\partial\over\partial\eta_{4}$}. For massless case, we can find a T1B-type reduction rule. Thus the only irreducible point corresponds the master integrals $(0,0,0,0,0)$. For the massive case, we can find T1B-type reduction rule for $\partial^{2}\over\partial\eta_{3}^{2}$}. Thus four points not reduced, i.e., (0,0,0,0,0),(1,0,0,0,0),(0,1,0,0,0),(0,0,1,0,0)(0,0,0,0,0),(1,0,0,0,0),(0,1,0,0,0),(0,0,1,0,0), correspond to the master integrals for massive sunset in the top sector.
V Example 2: the top sector of nonplanar double box diagrams
This is a nontrivial example. A nonplanar two-loop integral can be reduced in an elegant way. We choose a complete set of Lorentz scalars as
where the last two are ISPs. The kinematics specified as $k_{i}^{2}=0$, $i=1,2,4$ and $k_{1}\cdot k_{2}=s/2$, $k_{2}\cdot k_{3}=t/2$ and $k_{1}\cdot k_{3}=-(s+t)/2$.
The first round computation: The initial seeds are the $10$ DE’s from fundamental IBP relations. From them, we can solve $2$ T1A-type reduction rules for degree-one operators. $6$ T1A-type reduction rules for degree-two operators. There are also two T2A-type reduction rules. Using these reduction rules, irreducible lattice points are
${\cal U}_{11}$ $=$ $(0,n_{2},0,n_{4},n_{5},n_{6},n_{7},0,0),$
${\cal U}_{12}$ $=$ $(0,0,0,0,n_{5},0,n_{7},0,n_{9}),$
${\cal U}_{2}$ $=$ $(0,0,0,0,0,0,0,n_{8},n_{9}).\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(17)
The second round computation: Since the indices for $n_{1},n_{3}$ have been fully reduced, consider the action ${\partial\over\partial\eta_{i}},i\neq 1,3$ on above $10$ reduction rules. Among $70$ equations, there are $16$ nontrivial DE’s. Among them, there are $12$ degree-two DE’s and four degree-one DE’s. We can solve $10$ T1A-type reduction rules for degree-two operators. A T2A-type reduction rule, and a T2B-reduction rule are obtained. From the degree-one DE’s, three T2B-type reduction rules can be solved. Now the irreducible points are $(0,0,0,n_{4},n_{5},0,0,0,0)$, $(0,0,0,0,0,0,0,n_{8},0)$ and $(0,0,0,0,0,0,0,0,n_{9})$.
The third round computation: Since ${\partial\over\partial\eta_{2}},{\partial\over\partial\eta_{6}},{\partial\over\partial\eta_{7}}$ have been fully reduced, we consider the action of ${\partial\over\partial\eta_{i}},i=4,5,8,9$ only. To generate fewer equations, we act only on three degree-one reduction rules found in previous round. Using these $12$ DE’s, we can solve $3$ T1B-type reduction rules for degree-2 operators ${\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{6}}$, $\partial^{2}\over\partial\eta_{4}^{2}$} and $\partial^{2}\over\partial\eta_{5}^{2}$}. Furthermore, we can solve $\partial\over\partial\eta_{9}$} , $\partial\over\partial\eta_{8}$}, and $\partial\over\partial\eta_{5}$}. Using these reduction rules, reduce $n_{5},n_{8},n_{9}$ are reduced to zero, and by the T1B-reduction rule of $\partial^{2}\over\partial\eta_{4}^{2}$}, $n_{4}$ to $0,1$ are reduced to zero. So thus on the top sector, all integrals are reduced to two master integrals.
VI Conclusion and Outlook
In this work, we introduced a novel systematic approach to derive complete symbolic reduction rules for multi-loop Feynman integrals with generating functions. By the DEs for these generating functions, we obtain recurrence relations that enable the efficient reduction without the exponential proliferation of IBP identities in traditional methods. In other words, the IBP relations are solved analytically. Through explicit examples, including the sunset diagrams (both massive and massless), two-loop double-box, and two-loop non-planar integrals, we demonstrate the algorithm’s effectiveness in complex cases with high-degree numerators and high-power propagators.
With the proof of concepts in the note, the most important thing is to automate the aforementioned algorithm. In a longer companion paper, more details of the algorithm will be presented and cutting-edge examples like the reduction of two-loop five-point massless and massive Feynman integrals for collider phenomenology.
The combination of the idea in the paper and other reduction methods, as well as the application on more general input other than Feynman integrals, like energy-energy correlator and cosmological correlator, are also worth pursuing.
Acknowledgements.
Supplemental Material
In the supplemental material, we provide details of the examples in this paper, as well as the two-loop four-point planar integral reduction example.
Appendix A The topsector of sunset
@@
where for simplicity $\mu_{4}=\mu_{5}=0$ have been implied. Among $2^{3}=8$ possible generating function, only following four sectors are nonzero
@@
The first round of computation: The initial seed for top-sector $G_{11100}=G_{111}$ is six degree-2 DE’s from fundamental IBP relations. Focus on the degree-2 operators in these equations, we have following coefficient matrix
$\left(\begin{array}[]{c| c | c |c |c |c |l}[1,4]&[2,4]&[3,4]&[1,5]&[2,5]&[3,5]&{\rm IBP}\\ \hline\cr&&&&&2&{d\over\partial\ell_{1}}\ell_{1}\\ \hline\cr-2&&-2&-2&&&{d\over\partial\ell_{1}}\ell_{2}\\ \hline\cr-2&&-2&&&-2&{d\over\partial\ell_{1}}K\\ \hline\cr&-2&&&-2&-2&{d\over\partial\ell_{1}}\ell_{1}\\ \hline\cr&&2&&&&{d\over\partial\ell_{2}}\ell_{2}\\ \hline\cr&&-2&&-2&-2&{d\over\partial\ell_{2}}K\end{array}\right)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(28)
where $a,b]\equiv{\partial\over\partial\eta_{a}}{\partial\over\partial\eta_{b}}$.
Doing the Gauss elimination, we can get the proper combinations of these six equations. After solving them, we get six T1A-type reduction rules for six degree-2 operators ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{j}}$, $i=4,5;j=1,2,3$, which are (where $f_{\pm\pm\pm}=\pm m_{1}^{2}\pm m_{2}^{2}\pm m_{3}^{2}-K^{2}$)
$\left\{2{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{3}}G_{111}\right\}=\left\{{2m_{1}^{2}\over s_{0}}{\partial\over\partial\eta_{1}}G_{111}+{f_{+-+}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}\right\}$
$-\left\{(D-3)G_{111}+\eta_{4}{\partial\over\partial\eta_{4}}G_{111}-2\eta_{1}{\partial\over\partial\eta_{1}}G_{111}\right$.